OpenStudy (anonymous):

Find an equation for the line perpendicular to the tangent to the curve y=x^3 -9x+1 at the point (3,1)

OpenStudy (anonymous):

Differentiate the equation to find dy/dx. dy/dx =3x^2 - 9 slope at (3,1) is dy/dx at x=3 and y=1 so, slope of required line=m= 3(3)^2 -9=27-9=18 now using slope point form equation of tangent is (y-1)=18(x-3) y-1=18x-54 18x-y-53=0

OpenStudy (anonymous):

Is 18x-y-53=0 the answer?

OpenStudy (anonymous):

no its the equation of tangent which I mistakenly derived. Slope of the line perpendicular to the tangent to the curve is given as m=-dx/dy=-1/(dy/dx)=-1/18 Now find the equation in slope point form as done above.. Got it??

OpenStudy (anonymous):

No. I'm sorry but what would be the answer?

OpenStudy (anonymous):

By slope point form (y-1)=-(1/18)(x-3) 18y-18=-x+3 x+18y-21=0

OpenStudy (anonymous):

so x+18y-27=0?

OpenStudy (anonymous):

x+18y-21=0

OpenStudy (anonymous):

oh, right. Thanks

OpenStudy (anonymous):

its a wrong answer :(